The hypotenuse of a right triangle measures 90 meters and one leg measures 72 meters find the area of the triangle
Answer by checkley77(12844)

A^2+B^2=C^2
72^2+B^2=90^2
5,184+B^2=8,100
B^2=8,100-5,184
B^2=2,916
B=SQRT2,916
B=54 ANS. FOR THE OTHER LEG.
PROOF:
72^2+54^2=80^2
5,184+2,916=8,100
8,100=8,100
AREA=BH/2
AREA=54*72/2=3,888/2=1,944 METERS ANS. FOR THE AREA.